OpenStudy (anonymous):

Can someone help me with 5 questions?

OpenStudy (anonymous):

I can try.

OpenStudy (anonymous):

Megan and Tina each have loans for $7000 with a 3-year loan term. Megan pays 6% interest and Tina pays 4% interest. If Megan’s monthly payment is $213 and Tina’s monthly payment is $207, how much more in interest will Megan pay than Tina? A. $213 B. $216 C. $452 D. $668

OpenStudy (anonymous):

I think its A.

OpenStudy (anonymous):

Let me work it give me a min . ^^

OpenStudy (anonymous):

A seems to be right.

OpenStudy (anonymous):

What is the total cost to repay a credit card loan with a balance of $25,000, an APR of 15%, and a minimum monthly payment of $330?

OpenStudy (anonymous):

I don't know the answer for this.

OpenStudy (anonymous):

I'm work it one moment,.

OpenStudy (anonymous):

I want you to get this so to cal. total cost to want to change the APR rate percentage to a decimal.

OpenStudy (anonymous):



OpenStudy (anonymous):

next you add that to 25k. and you get 28750

OpenStudy (anonymous):

Josh and Michael each have a loan for $6200 with a 4% interest rate. Josh has a monthly payment of $183, and Michael has a monthly payment of $114. How much more in interest will Michael pay if his loan term is 5 years and Josh’s loan term is 3 years? A. $114 B. $183 C. $252 D. $388

OpenStudy (anonymous):

I think its A.

OpenStudy (anonymous):

Or B

OpenStudy (anonymous):

A appears to be right.

OpenStudy (anonymous):

On an easy-access loan for $1500, what is the total cost to repay the loan in 30 days with a 22% monthly fee?

OpenStudy (anonymous):

i don't know this one either

OpenStudy (anonymous):

@amistre64 I will ask him to double check everything and help you further, I have to go >< sorry

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem and Great luck in your studies ^^.

OpenStudy (amistre64):

total cost = total payment = loan + interest therefore interest = total payments - loan 36(p1) - L = int1 36(p2) - L = int2 int1 - int2 is the difference in interest: (36(p1)- L) - (36(p2)- L) 36(p1) - 36(p2) 36(p1-p2) the difference in interest is simply the difference in payment over the 3 years.

OpenStudy (anonymous):

what is the answer though?

OpenStudy (amistre64):

im not going to add and multiply for you ....

OpenStudy (anonymous):

what would the p stand for?

OpenStudy (amistre64):

payment, what is the difference in their monthly payments?

OpenStudy (anonymous):

i dont get it

OpenStudy (amistre64):

you dont get what? the difference in the payments? this is starting at your first question. what are the payments?

OpenStudy (amistre64):

question 2 i disagree with as well. assume you have a bank loan of 25000 and are making payments of 330 a month, how long dies it take to pay off the loan? Bk^n = P(k^n-1)/(k-1) B(k-1) k^n = Pk^n -P P = Pk^n -B(k-1) k^n P = k^n(P -B(k-1)) P/(P -B(k-1)) = k^n log(P/(P -B(k-1))) = n log(k) log(P/(P -B(k-1)))/log(k) = n

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=log%28330%2F%28330+-25000%28%281%2B.15%2F12%29-1%29%29%29%2Flog%28%281%2B.15%2F12%29%29 itll take about 236 payments to pay off. and a tad more

OpenStudy (anonymous):

is that one for question 2?

OpenStudy (amistre64):

thats what i worked out for 2 yes. how long is 236 months?

OpenStudy (anonymous):

About 19 years

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=25000k%5E%28236%29-330%28k%5E%28236%29-1%29%2F%28k-1%29%2C+k%3D1%2B.15%2F12 see, when n= 236 we have just a tad bit left to go, and if n=237 we are negative ... so the last payment is the compounding of the 236th

OpenStudy (amistre64):

yeah, long time :)

OpenStudy (amistre64):

236 payments, plus 136.32(1+.15/12) should do it

OpenStudy (amistre64):

about 78 018, which is a far cry from 28 750

OpenStudy (amistre64):

all that was to determine how long it would take to payoff the amount. I tend to use my own formula that makes much more sense to me.

OpenStudy (anonymous):

are we still on 2

OpenStudy (amistre64):

yes, i havent indicated otherwise.

OpenStudy (amistre64):

Josh and Michael each have a loan for $6200 with a 4% interest rate. Josh has a monthly payment of $183, and Michael has a monthly payment of $114. How much more in interest will Michael pay if his loan term is 5 years and Josh’s loan term is 3 years? this is like the first, except for different time spans as well 5*12*p1 - L = int1 3*12*p2 - L = int2 int1 - int2 = 5*12*p1 - 3*12*p2 12(5p1 - 3p2)

OpenStudy (amistre64):

i do not get A ... therefore i cant agree with any of the solutions that were presented

OpenStudy (amistre64):

so, where do you want to start the study with?

OpenStudy (anonymous):

I finished.. Thank you very much!!!!

OpenStudy (amistre64):

ok, did you modify your choices any? or just went with what you had?

OpenStudy (anonymous):

i modified

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

1) 207 and 213 a difference of 6; 36(6) is 216 2) finding n was critical, the nP + the compounded remains is about 78018 3) 60 payments of 114 - 36 payments of 183 60(114) - 36(183) = 252 those would have been my choices.